# Understanding and Leveraging the Expert Specialization of Context Faithfulness in Mixture-of-Experts LLMs

### v1

@@ Abstract
Jun Bai1 Minghao Tong1,2 Yang Liu1 Zixia Jia1,🖂 Zilong Zheng1,🖂 1 State Key Laboratory of General Artificial Intellligence, BIGAI 2 School of Computer Science, Wuhan University {baijun, tongminghao1, liuyang, jiazixia, zlzheng}@bigai.ai
@@ 1 Introduction
Faithfulness to the provided context is essential for ensuring the reliability and coherence of responses in many context-dependent scenarios, such as long sequence processing Li et al. (2024); Wu et al. (2025), In-Context Learning (ICL) Zhang et al. (2025); Qi et al. (2025), and Retrieval-Augmented Generation (RAG) Shi et al. (2024); Sun et al. (2025). Despite their remarkable fluency, Large Language Models (LLMs) often generate outputs that are only loosely grounded in the given context or, in more concerning instances, hallucinate information not supported by it Zhou et al. (2023); Chuang et al. (2024); Huang et al. (2025).
Figure 1: A case from NQ-Swap Longpre et al. (2021) where MoE experts exhibit different tendencies of context faithfulness (answer is underlined in green).
Prior work has proposed sophisticated prompting Zhou et al. (2023), decoding Shi et al. (2024), and alignment Bi et al. (2024) techniques to enhance context faithfulness—the ability to accurately attend to and integrate relevant contextual information during response generation. More recently, expert specialization in Mixture-of-Experts (MoE) models has emerged as a promising direction Wang et al. (2024b), potentially enabling more targeted optimization of model capacity for context utilization. Previous studies have shown that MoE experts tend to specialize in processing different aspects of input. The router network typically activates distinct experts when handling tokens from diverse tasks Jiang et al. (2024), domains Muennighoff et al. (2024), and syntactic units Antoine et al. (2025). Building on these observations, we identify that experts in MoE models exhibit varying degrees of context faithfulness, as illustrated in Figure 1. This raises an intriguing question: Do context-faithful experts exist within MoE models?
In this work, we present a systematic analysis of expert specialization in MoE LLMs, with a particular focus on their ability to leverage contextual information. A common approach to identifying experts responsible for specific functionalities involves analyzing expert activation frequencies across layers Jiang et al. (2024); Wang et al. (2024b), under the assumption that the most frequently activated experts are those most relevant to the task. However, this assumption is undermined by the load balancing constraint imposed during pretraining Dai et al. (2024), which enforces uniform expert usage and consequently limits the router network’s ability to select the most beneficial experts Dai et al. (2022). As a result, experts identified through standard activation-based heuristics may not accurately reflect optimal specialization for context-dependent behavior.
@@ 2.1 Context-dependent Tasks
In context-dependent task, the correct prediction or output depends not only on the input query $q$ but also on an additional context $c$, which provides essential supporting information Kazi et al. (2023); Fan et al. (2024); Bi et al. (2024). Formally, let $\mathcal{Q}$ be the space of queries, $\mathcal{C}$ be the space of possible contexts, and $\mathcal{Y}$ be the space of outputs. A context-dependent task is defined by a function:
@@ 2.2 Mixture-of-Experts LLMs
MoE is an architecture where the Feed-Forward Network (FFN) Vaswani et al. (2017) are replaced by MoE modules to efficiently increase model capacity Xue et al. (2024). Each MoE layer consists of $N_{e}$ parallel experts $e$ (sharing FFN structure). For each input token, a subset of $k$ experts is activated (i.e. top-$k$) based on learned gating scores computed by a router network Dai et al. (2022).
@@ 3.1 Router Lens
To identify experts relevant to a specific capability, prior work often uses expert activation frequency as a proxy—experts that are activated more frequently are assumed to be more important Jiang et al. (2024); Wang et al. (2024b). However, in pretrained MoE models, the router network is typically optimized with a strong load balancing constraint to enforce uniform expert usage Wang et al. (2024a). While this constraint improves training stability and computational efficiency, it can obscure the natural emergence of expert specialization for particular capabilities such as context utilization. As a result, models may struggle to accurately identify and leverage the most beneficial experts for context-dependent tasks.
@@ 3.2 Experimental Setting
We evaluate our approach on several widely used datasets for context-dependent tasks, including SQuAD Rajpurkar et al. (2016), NQ Kwiatkowski et al. (2019), HotpotQA Yang et al. (2018), NQ-Swap Longpre et al. (2021), and ConfiQA (Multi-Conflicts subset) Bi et al. (2024). Dataset statistics are summarized in Table 6. Among these, SQuAD, NQ, and HotpotQA are classic question answering benchmarks that span a spectrum of reasoning challenges upon context, from local comprehension to multi-hop reasoning. Moreover, we also include two counterfactual QA benchmarks: NQ-Swap and ConfiQA, offering a more challenging and complementary evaluation of a model’s ability to identify and rely on the correct evidence. For all datasets, we report Exact Match (EM) and token-level F1 scores as evaluation metrics. For additional dataset details, please refer to Appendix B.
@@
We select four widely used open-source MoE-based LLMs for evaluation: OLMoE-1B-7B Muennighoff et al. (2024), DeepSeek-V2-Lite DeepSeek-AI et al. (2024), MiniCPM-MoE-8x2B Hu et al. (2024), and Mixtral-8x7B Jiang et al. (2024). These models cover a diverse range of configurations in terms of the number of experts (from 8 to 64) and overall model sizes (from 7B to 47B), providing a comprehensive testbed for analyzing expert behavior in context-dependent scenarios. Detailed configurations for each model are listed in Table 7. To ensure consistent and deterministic outputs across models, we adopt greedy decoding Germann (2003) for all experiments. For additional model details, please refer to Appendix C.
@@ 3.3 Results of Expert Elicitation
To assess whether the performance gains from the newly selected experts arise solely from enhanced compositional reasoning rather than effective context utilization, we also conduct experiments on the CounterFact dataset Meng et al. (2022), where the model must rely on the provided counterfactual context to answer a simple question correctly. The task does not require complex reasoning, making it well-suited for isolating and evaluating the role of context-faithful experts. Table 2 reports the EM and F1 scores of OLMoE-1B-7B and MiniCPM-MoE-8x2B on CounterFact. We observe substantial performance improvements after router tuning, further providing strong evidence for the existence of context-faithful experts.
@@ 3.4 Consistency of Context-Faithful Experts
We further investigate whether the same context-faithful experts are consistently activated across different context-dependent tasks. To this end, we analyze expert activation patterns in OLMoE-1B-7B and MiniCPM-MoE-8×2B. For each input sample, we calculate the activation frequency of context-faithful experts across all layers and concatenate these values into a feature vector. We then apply t-SNE Van der Maaten and Hinton (2008) to project these vectors into a 2D space for visualization. As shown in Figure 4, samples from different datasets form clearly separable clusters in both OLMoE-1B-7B and MiniCPM-MoE-8×2B. This indicates that the router learns task-specific activation patterns for context-faithful experts, demonstrating that the model adapts its routing behavior based on the contextual requirements of each task, highlighting both the interpretability and task-awareness enabled by the router-tuning mechanism.
@@ 3.5 Inner Working Mechanism of Context-faithful Experts
Moreover, we investigate how context-faithful experts contribute to improving context faithfulness. In Transformer-based models, the self-attention mechanism plays a central role in perceiving and integrating contextual information Sun et al. (2025). By assigning higher attention scores to relevant context tokens, the model can more effectively utilize the provided context. To assess whether context-faithful experts enhance this mechanism, we examine whether their activation leads to increased attention over contextual tokens compared to the untuned model. To this end, we introduce the Context Attention Gain (CAG) metric.
@@
$\text{CAG}^{(\ell)}=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\left(\frac{\alpha_{i,\text{rt}}^{(\ell)}-\alpha_{i,\text{base}}^{(\ell)}}{\alpha_{i,\text{base}}^{(\ell)}}\right)$ (11)
@@
In addition, we define Answer Probability Gain (APG) metric to examine how context-faithful experts influence the model’s intermediate decision-making process. Let $\mathbf{h}_{i,last}^{(\ell)}\in\mathbb{R}^{d}$ be the hidden state output by the MoE module at layer $\ell$ for the last token of input $i$ where $d$ is hidden size, and let $y_{i}$ be its true answer tokens. We employ Logit Lens Dar et al. (2023), using the projection matrix $\mathbf{W}\in\mathbb{R}^{V\times d}$ (where $V$ is the vocabulary size) of language model head, to compute the probability $p_{m}^{(\ell)}(y_{i})$ assigned to $y_{i}$ at layer $\ell$ under model $m\in\{\mathrm{base},\mathrm{rt}\}$ where $\mathrm{rt}$ refers to “router-tuned”. We then compute the average probability gain of the correct answer over all $N_{s}$ samples:
$\text{APG}^{(\ell)}=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\Bigl{[}\frac{p_{\mathrm{rt}}^{(\ell)}(y_{i})-p_{\mathrm{base}}^{(\ell)}(y_{i})}{p_{\mathrm{base}}^{(\ell)}(y_{i})}\Bigr{]}$ (12)
@@ 4.1 Context-faithful Expert Fine-Tuning
Table 3: Performance comparison of Fully Fine-Tuning (FFT), Expert-Specialized Fine-Tuning (ESFT) (Wang et al., 2024b), and Context-faithful Expert Fine-Tuning (CEFT). Bold numbers indicate the best performance, while underlined numbers denote the second-best.
@@ 4.2 Empirical Results
To evaluate the effectiveness of the proposed CEFT approach, we compare it against two baseline strategies: standard Fully Fine-Tuning (FFT) and Expert-Specialized Fine-Tuning (ESFT) Wang et al. (2024b). While ESFT also targets expert-level adaptation by only tuning a subset of experts deemed relevant to the task, it relies on the original, untuned router network to identify these experts, which may result in sub-optimal expert selection due to the influence of load balancing constraint and the limited task-awareness of the untuned router. In contrast, CEFT first adapts the router to reveal truly context-faithful experts—those that significantly enhance context-dependent reasoning—and then fine-tunes only these high-impact experts for more precise and efficient adaptation. For implementation details, please refer to the Appendix A. As shown in Table 3, CEFT consistently matches or surpasses FFT and ESFT across MoE models and benchmarks, indicating its parameter-effectiveness in leveraging contextual information for improving generalization in context-dependent tasks.
@@
We further evaluate OLMoE-1B-7B and MiniCPM-MoE-8x2B on MMLU Hendrycks et al. (2021) to assess the impact of RT, FFT, and CEFT on the models’ original capabilities. As shown in Table 4, the decline in MMLU performance is approximately proportional to the number of trainable parameters, with CEFT being substantially less susceptible to catastrophic forgetting compared to FFT.
@@ 5 Related works
Context faithfulness refers to the extent to which a model’s output remains accurate, consistent, and grounded in the provided context. Unlike factual consistency, it emphasizes alignment specifically with the input context rather than with external world knowledge Zhou et al. (2023); Xu et al. (2024). In tasks such as RAG, maintaining context faithfulness is critical to ensure the reliability and trustworthiness of generated content Zhou et al. (2024). Recent studies have shown that LLMs often generate responses that are fluent yet contextually unfaithful Du et al. (2024); Xie et al. (2024). Various techniques, including prompt engineering Zhou et al. (2023), decoding constraints Shi et al. (2024), attention amplification Sun et al. (2025), mechanistic approaches Yu et al. (2023); Ortu et al. (2024); Minder et al. (2025), and learning-based methods such as RAG-fashioned fine-tuning Zhang et al. (2024) and reinforcement learning with context-aware rewards Bi et al. (2024), have been proposed to improve grounding in the given context. Differently, our work focuses on an under-explored direction: leveraging the expert specialization in terms of the context utilization to improve the MoE performance on context-dependent tasks.
@@
The MoE architecture replaces the standard FFN layer Vaswani et al. (2017) with a modular MoE layer comprising multiple parallel experts and a router network to sparsely activate top-$k$ experts Xue et al. (2024). During the pretraining stage of MoE, a load balancing loss is typically incorporated to force balanced expert utilization Dai et al. (2024), which encourages the formation of expert specialization. Recently, a lot of work has focused on the understanding of expert specialization, finding that different experts are responsible for different tokens Muennighoff et al. (2024), domains Jiang et al. (2024), tasks Wang et al. (2024b), semantic units Li and Zhou (2025), or syntactic units Antoine et al. (2025). These findings facilitate the targeted optimization of the specialized experts Wang et al. (2024b). Inspired by these, our work investigates the expert specialization of context faithfulness.

### v3

@@ Abstract
Jun Bai1  Minghao Tong1,2  Yang Liu1  Zixia Jia1,🖂  Zilong Zheng1,🖂 1 State Key Laboratory of General Artificial Intelligence, BIGAI 2 School of Computer Science, Wuhan University {baijun, liuyang, jiazixia, zlzheng}@bigai.ai, tongminghao@whu.edu.cn
@@ 1 Introduction
Faithfulness to the provided context is essential for ensuring the reliability and coherence of responses in many context-dependent scenarios, such as long sequence processing DBLP:conf/acl/LiWZZ24; DBLP:journals/corr/abs-2502-18890, In-Context Learning (ICL) DBLP:conf/acl/0003LBLWLR25; DBLP:conf/iclr/QiYJ0LZ0Z25, and Retrieval-Augmented Generation (RAG) DBLP:conf/naacl/ShiHLTZY24; DBLP:conf/iclr/SunZ0XZYSL25. Despite their remarkable fluency, Large Language Models (LLMs) often generate outputs that are only loosely grounded in the given context or, in more concerning instances, hallucinate information not supported by it DBLP:conf/emnlp/ZhouZPC23; DBLP:conf/emnlp/ChuangQHKKG24; DBLP:journals/tois/HuangYMZFWCPFQL25.
Figure 1: A case from NQ-Swap DBLP:conf/emnlp/LongprePCRD021/nq-swap where MoE experts exhibit different tendencies of context faithfulness (answer is underlined in red).
Prior work has proposed sophisticated prompting DBLP:conf/emnlp/ZhouZPC23, decoding DBLP:conf/naacl/ShiHLTZY24, and alignment DBLP:journals/corr/abs-2412-15280/confliqa techniques to enhance context faithfulness—the ability to accurately attend to and integrate relevant contextual information during response generation. More recently, expert specialization in Mixture-of-Experts (MoE) models has emerged as a promising direction DBLP:conf/emnlp/WangCDXLW24/esft, potentially enabling more targeted optimization of model capacity for context utilization. Previous studies have shown that MoE experts tend to specialize in processing different aspects of input. The router network typically activates distinct experts when handling tokens from diverse tasks DBLP:journals/corr/abs-2401-04088/mixtral, domains DBLP:journals/corr/abs-2409-02060/olmoe, and syntactic units DBLP:conf/coling/AntoineBL25. Building on these observations, we identify that experts in MoE models exhibit varying degrees of context faithfulness, as illustrated in Figure 1. This raises an intriguing question: Do context-faithful experts exist within MoE models?
In this work, we present a systematic analysis of expert specialization in MoE LLMs, with a particular focus on their ability to leverage contextual information. A common approach to identifying experts responsible for specific functionalities involves analyzing expert activation frequencies across layers DBLP:journals/corr/abs-2401-04088/mixtral; DBLP:conf/emnlp/WangCDXLW24/esft, under the assumption that the most frequently activated experts are those most relevant to the task. However, this assumption is undermined by the load balancing constraint imposed during pretraining DBLP:conf/acl/DaiDZXGCLZYWXLH24/DeepSeekMOE, which enforces uniform expert usage and consequently limits the router network’s ability to select the most beneficial experts DBLP:conf/acl/Dai0MZSCW22/router. As a result, experts identified through standard activation-based heuristics may not accurately reflect optimal specialization for context-dependent behavior.
@@ 2.1 Context-dependent Tasks
In context-dependent task, the correct prediction or output depends not only on the input query $q$ but also on an additional context $c$, which provides essential supporting information DBLP:journals/air/KaziKD23; DBLP:conf/kdd/FanDNWLYCL24; DBLP:journals/corr/abs-2412-15280/confliqa. Formally, let $\mathcal{Q}$ be the space of queries, $\mathcal{C}$ be the space of possible contexts, and $\mathcal{Y}$ be the space of outputs. A context-dependent task is defined by a function:
@@ 2.2 Mixture-of-Experts LLMs
MoE is an architecture where the Feed-Forward Network (FFN) DBLP:conf/nips/VaswaniSPUJGKP17/transformer_first are replaced by MoE modules to efficiently increase model capacity DBLP:conf/icml/XueZFNZZ024/openmoe. Each MoE layer consists of $N_{e}$ parallel experts $e$ (sharing FFN structure). For each input token, a subset of $k$ experts is activated (i.e. top-$k$) based on learned gating scores computed by a router network DBLP:conf/acl/Dai0MZSCW22/router.
@@ 3.1 Router Lens
To identify experts relevant to a specific capability, prior work often uses expert activation frequency as a proxy—experts that are activated more frequently are assumed to be more important DBLP:journals/corr/abs-2401-04088/mixtral; DBLP:conf/emnlp/WangCDXLW24/esft. However, in pretrained MoE models, the router network is typically optimized with a strong load balancing constraint to enforce uniform expert usage DBLP:journals/corr/abs-2408-15664. While this constraint improves training stability and computational efficiency, it can obscure the natural emergence of expert specialization for particular capabilities such as context utilization. As a result, models may struggle to accurately identify and leverage the most beneficial experts for context-dependent tasks.
@@ 3.2 Experimental Setting
We evaluate our approach on several widely used datasets for context-dependent tasks, including SQuAD DBLP:conf/emnlp/RajpurkarZLL16/squad, NQ DBLP:journals/tacl/KwiatkowskiPRCP19/nq, HotpotQA DBLP:conf/emnlp/Yang0ZBCSM18/hotpotqa, NQ-Swap DBLP:conf/emnlp/LongprePCRD021/nq-swap, and ConfiQA (Multi-Conflicts subset) DBLP:journals/corr/abs-2412-15280/confliqa. Dataset statistics are summarized in Table 6. Among these, SQuAD, NQ, and HotpotQA are classic question answering benchmarks that span a spectrum of reasoning challenges upon context, from local comprehension to multi-hop reasoning. Moreover, we also include two counterfactual QA benchmarks: NQ-Swap and ConfiQA, offering a more challenging and complementary evaluation of a model’s ability to identify and rely on the correct evidence. For all datasets, we report Exact Match (EM) and token-level F1 scores as evaluation metrics. For additional dataset details, please refer to Appendix B.
@@
We select four widely used open-source MoE-based LLMs for evaluation: OLMoE-1B-7B DBLP:journals/corr/abs-2409-02060/olmoe, DeepSeek-V2-Lite DBLP:journals/corr/abs-2405-04434/deepseek-v2, MiniCPM-MoE-8x2B DBLP:journals/corr/abs-2404-06395/minicpm, and Mixtral-8x7B DBLP:journals/corr/abs-2401-04088/mixtral. These models cover a diverse range of configurations in terms of the number of experts (from 8 to 64) and overall model sizes (from 7B to 47B), providing a comprehensive testbed for analyzing expert behavior in context-dependent scenarios. Detailed configurations for each model are listed in Table 7. To ensure consistent and deterministic outputs across models, we adopt greedy decoding DBLP:conf/naacl/Germann03 for all experiments. For additional model details, please refer to Appendix C.
@@ 3.3 Results of Expert Elicitation
To assess whether the performance gains from the newly selected experts arise solely from enhanced compositional reasoning rather than effective context utilization, we also conduct experiments on the CounterFact dataset DBLP:conf/nips/MengBAB22, where the model must rely on the provided counterfactual context to answer a simple question correctly. The task does not require complex reasoning, making it well-suited for isolating and evaluating the role of context-faithful experts. Table 2 reports the EM and F1 scores of OLMoE-1B-7B and MiniCPM-MoE-8x2B on CounterFact. We observe substantial performance improvements after router tuning, further providing strong evidence for the existence of context-faithful experts.
@@ 3.4 Consistency of Context-Faithful Experts
We further investigate whether the same context-faithful experts are consistently activated across different context-dependent tasks. To this end, we analyze expert activation patterns in OLMoE-1B-7B and MiniCPM-MoE-8×2B. For each input sample, we calculate the activation frequency of context-faithful experts across all layers and concatenate these values into a feature vector. We then apply t-SNE van2008visualizing to project these vectors into a 2D space for visualization. As shown in Figure 4, samples from different datasets form clearly separable clusters in both OLMoE-1B-7B and MiniCPM-MoE-8×2B. This indicates that the router learns task-specific activation patterns for context-faithful experts, demonstrating that the model adapts its routing behavior based on the contextual requirements of each task, highlighting both the interpretability and task-awareness enabled by the router-tuning mechanism.
@@ 3.5 Inner Working Mechanism of Context-faithful Experts
Moreover, we investigate how context-faithful experts contribute to improving context faithfulness. In Transformer-based models, the self-attention mechanism plays a central role in perceiving and integrating contextual information DBLP:conf/iclr/SunZ0XZYSL25. By assigning higher attention scores to relevant context tokens, the model can more effectively utilize the provided context. To assess whether context-faithful experts enhance this mechanism, we examine whether their activation leads to increased attention over contextual tokens compared to the untuned model. To this end, we introduce the Context Attention Gain (CAG) metric.
@@
$\text{CAG}^{(\ell)}=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\frac{\alpha_{i,\text{rt}}^{(\ell)}-\alpha_{i,\text{base}}^{(\ell)}}{\alpha_{i,\text{base}}^{(\ell)}}$ (11)
@@
In addition, we define Answer Probability Gain (APG) metric to examine how context-faithful experts influence the model’s intermediate decision-making process. Let $\mathbf{h}_{i,last}^{(\ell)}\in\mathbb{R}^{d}$ be the hidden state output by the MoE module at layer $\ell$ for the last token of input $i$ where $d$ is hidden size, and let $y_{i}$ be its true answer tokens. We employ Logit Lens DBLP:conf/acl/DarG0B23/logitlens, using the projection matrix $\mathbf{W}\in\mathbb{R}^{V\times d}$ (where $V$ is the vocabulary size) of language model head, to compute the probability $p_{m}^{(\ell)}(y_{i})$ assigned to $y_{i}$ at layer $\ell$ under model $m\in\{\mathrm{base},\mathrm{rt}\}$ where $\mathrm{rt}$ refers to “router-tuned”. We then compute the average probability gain of the correct answer over all $N_{s}$ samples:
$\text{APG}^{(\ell)}=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\frac{p_{\mathrm{rt}}^{(\ell)}(y_{i})-p_{\mathrm{base}}^{(\ell)}(y_{i})}{p_{\mathrm{base}}^{(\ell)}(y_{i})}$ (12)
@@ 4.1 Context-faithful Expert Fine-Tuning
Table 3: Performance comparison of Fully Fine-Tuning (FFT), Expert-Specialized Fine-Tuning (ESFT) (DBLP:conf/emnlp/WangCDXLW24/esft), and Context-faithful Expert Fine-Tuning (CEFT). Bold numbers indicate the best performance, while underlined numbers denote the second-best.
@@ 4.2 Empirical Results
To evaluate the effectiveness of the proposed CEFT approach, we compare it against two baseline strategies: standard Fully Fine-Tuning (FFT) and Expert-Specialized Fine-Tuning (ESFT) DBLP:conf/emnlp/WangCDXLW24/esft. While ESFT also targets expert-level adaptation by only tuning a subset of experts deemed relevant to the task, it relies on the original, untuned router network to identify these experts, which may result in sub-optimal expert selection due to the influence of load balancing constraint and the limited task-awareness of the untuned router. In contrast, CEFT first adapts the router to reveal truly context-faithful experts—those that significantly enhance context-dependent reasoning—and then fine-tunes only these high-impact experts for more precise and efficient adaptation. For implementation details, please refer to the Appendix A. As shown in Table 3, CEFT consistently matches or surpasses FFT and ESFT across MoE models and benchmarks, indicating its parameter-effectiveness in leveraging contextual information for improving generalization in context-dependent tasks.
@@
We further evaluate OLMoE-1B-7B and MiniCPM-MoE-8x2B on MMLU DBLP:conf/iclr/HendrycksBBZMSS21 to assess the impact of RT, FFT, and CEFT on the models’ original capabilities. As shown in Table 4, the decline in MMLU performance is approximately proportional to the number of trainable parameters, with CEFT being substantially less susceptible to catastrophic forgetting compared to FFT.
@@ 5 Related works
Context faithfulness refers to the extent to which a model’s output remains accurate, consistent, and grounded in the provided context. Unlike factual consistency, it emphasizes alignment specifically with the input context rather than with external world knowledge DBLP:conf/emnlp/ZhouZPC23; DBLP:conf/emnlp/XuQGW0ZX24. In tasks such as RAG, maintaining context faithfulness is critical to ensure the reliability and trustworthiness of generated content DBLP:journals/corr/abs-2409-10102. Recent studies have shown that LLMs often generate responses that are fluent yet contextually unfaithful DBLP:conf/acl/DuSSWSC24; DBLP:conf/iclr/Xie0CL024. Various techniques, including prompt engineering DBLP:conf/emnlp/ZhouZPC23, decoding constraints DBLP:conf/naacl/ShiHLTZY24, attention amplification DBLP:conf/iclr/SunZ0XZYSL25, mechanistic approaches DBLP:conf/emnlp/YuMP23; DBLP:conf/acl/OrtuJDSCS24; DBLP:conf/iclr/MinderDSMW0C25, and learning-based methods such as RAG-fashioned fine-tuning DBLP:journals/corr/abs-2403-10131 and reinforcement learning with context-aware rewards DBLP:journals/corr/abs-2412-15280/confliqa, have been proposed to improve grounding in the given context. Differently, our work focuses on an under-explored direction: leveraging the expert specialization in terms of the context utilization to improve the MoE performance on context-dependent tasks.
@@
The MoE architecture replaces the standard FFN layer DBLP:conf/nips/VaswaniSPUJGKP17/transformer_first with a modular MoE layer comprising multiple parallel experts and a router network to sparsely activate top-$k$ experts DBLP:conf/icml/XueZFNZZ024/openmoe. During the pretraining stage of MoE, a load balancing loss is typically incorporated to force balanced expert utilization DBLP:conf/acl/DaiDZXGCLZYWXLH24/DeepSeekMOE, which encourages the formation of expert specialization. Recently, a lot of work has focused on the understanding of expert specialization, finding that different experts are responsible for different tokens DBLP:journals/corr/abs-2409-02060/olmoe, domains DBLP:journals/corr/abs-2401-04088/mixtral, tasks DBLP:conf/emnlp/WangCDXLW24/esft, semantic units DBLP:conf/iclr/LiZ25, or syntactic units DBLP:conf/coling/AntoineBL25. These findings facilitate the targeted optimization of the specialized experts DBLP:conf/emnlp/WangCDXLW24/esft. Inspired by these, our work investigates the expert specialization of context faithfulness.